# A rigidity theorem for complex Kleinian groups

### v1

@@ 4. Simplicity of the Zariski closure
Suppose $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a strongly irreducible $\mathsf{P}_{k}$-Anosov subgroup and when $k>1$ assume that $\Gamma$ is also $\operatorname{\mathsf{P}}_{k-1}$-Anosov. If $\Lambda_{k}(\Gamma)$ is a ${\mathcal{C}}^{1}$-submanifold of $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$, then the identity component of the real Zariski closure of $\Gamma$ is a simple Lie group with trivial center.
Example 4.2.
When $k>1$, there exist strongly irreducible $\mathsf{P}_{k}$-Anosov subgroups where the limit set is a ${\mathcal{C}}^{1}$-submanifold of $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{C}}^{d})$, but the real Zariski closure of $\Gamma$ is not simple. For instance, suppose $\Gamma_{0}\subset\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ is a uniform lattice, and let $\Gamma_{0}^{\prime}\subset\operatorname{\mathsf{SL}}(2,\operatorname{\mathbb{C}})$ denotes its lift to $\operatorname{\mathsf{SL}}(2,\operatorname{\mathbb{C}})$. By replacing $\Gamma_{0}^{\prime}$ by a finite index subgroup we can assume that $\Gamma_{0}^{\prime}$ admits a surjective homomorphism to a non-abelian free group, see Agol–Groves–Manning [1]. Then, since non-abelian free groups admit Zariski dense representations into $\mathsf{SU}(k)$, we can construct a Zariski dense representation $\tau_{0}^{\prime}:\Gamma_{0}\rightarrow\mathsf{SU}(k)$.
Next let $\tau^{\prime}:\Gamma_{0}^{\prime}\rightarrow\operatorname{\mathsf{SL}}(\operatorname{\mathbb{C}}^{2}\otimes\operatorname{\mathbb{C}}^{k})$ be the representation defined by
$\tau(\gamma)(v\otimes w)=(\gamma(v))\otimes(\tau_{0}(\gamma)(w)),$
and let $\tau:=\pi\circ\tau^{\prime}$, where $\pi:\operatorname{\mathsf{SL}}(\operatorname{\mathbb{C}}^{2}\otimes\operatorname{\mathbb{C}}^{k})\to\mathsf{PSL}(\operatorname{\mathbb{C}}^{2}\otimes\operatorname{\mathbb{C}}^{k})$ is the quotient map. Using the fact that $\Gamma_{0}<\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ is $\operatorname{\mathsf{P}}_{1}$-Anosov and has limit set $\Lambda_{1}(\Gamma)=\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{2})$, one can check that $\Gamma:=\tau(\Gamma_{0}^{\prime})$ is $\mathsf{P}_{k}$-Anosov with limit set
$\Lambda_{k}(\Gamma)=\{L\otimes\operatorname{\mathbb{C}}^{k}:L\in\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{2})\}.$
Then $\Lambda_{k}(\Gamma)$ is a ${\mathcal{C}}^{1}$-submanifold and the real Zariski closure of $\Gamma$ is isomorphic to
$\mathsf{P}(\operatorname{\mathsf{SL}}(2,\operatorname{\mathbb{C}})\oplus\mathsf{SU}(k)).$
4.1. Anosov groups in general semisimple Lie groups
A Lie group is reductive if its Lie algebra decomposes as a direct sum of simple Lie algebras and an Abelian Lie algebra (which is necessarily the center of its Lie algebra). Equivalently, it admits a representation that is completely reducible, i.e. decomposes into a direct sum of irreducible representations. A reductive Lie group is semisimple if its Lie algebra has trivial center.
If a Lie group $\operatorname{\mathsf{G}}$ is semisimple, then its Lie algebra $\mathfrak{g}$ admits a Cartan decomposition, i.e. we may write
$\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p},$
where $\mathfrak{k}\subset\mathfrak{g}$ is a maximal compact Lie subalgebra and $\mathfrak{p}$ is the orthogonal complement of $\mathfrak{k}$ with respect to the Killing form. Let $\mathfrak{a}_{\operatorname{\mathsf{G}}}\subset\mathfrak{p}$ be a maximal Abelian subspace. For every linear map $\alpha\in\mathfrak{a}_{\operatorname{\mathsf{G}}}^{*}$, let
$\mathfrak{g}_{\alpha}:=\{X\in\mathfrak{a}_{\operatorname{\mathsf{G}}}:[X,Y]=\alpha(X)Y\text{ for all }Y\in\mathfrak{a}_{\operatorname{\mathsf{G}}}\}.$
We say that $\alpha\in\mathfrak{a}_{\operatorname{\mathsf{G}}}^{*}$ is a restricted root of $\operatorname{\mathsf{G}}$ if $\mathfrak{g}_{\alpha}$ is non-zero. Let $\Phi$ denote the set of restricted simple roots of $\operatorname{\mathsf{G}}$. Then we have the restricted root space decomposition
$\mathfrak{g}=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Phi}\mathfrak{g}_{\alpha}.$
Choose a vector $X_{0}\in\mathfrak{a}_{\operatorname{\mathsf{G}}}$ such that $\alpha(X_{0})\neq 0$ for all $\alpha\in\Phi$. Then we may decompose $\Phi=\Phi_{-}\cup\Phi_{+}$, where $\Phi_{\pm}:=\{\alpha\in\Phi:{\rm sgn}(\alpha(X_{0}))=\pm 1\}$. It is a fact that $\Phi_{-}=-\Phi_{+}$, and that there is a basis $\Delta_{\operatorname{\mathsf{G}}}\subset\Phi_{+}$ of $\mathfrak{a}_{\operatorname{\mathsf{G}}}^{*}$ such that every linear map in $\Phi^{+}$ can be written as a non-negative linear combination of the elements in $\Delta_{\operatorname{\mathsf{G}}}$. The linear maps in $\Delta_{\operatorname{\mathsf{G}}}$ are called the restricted simple roots of $\operatorname{\mathsf{G}}$.
Given a non-empty subset $\theta\subset\Delta_{\operatorname{\mathsf{G}}}$, let
$\mathfrak{u}_{\theta}:=\bigoplus_{\alpha\in\Phi\setminus\operatorname{Span}_{\operatorname{\mathbb{R}}}(\Delta_{\operatorname{\mathsf{G}}}\setminus\theta)}\mathfrak{g}_{\alpha}.$
The $\theta$-standard parabolic subgroup is the subgroup $\operatorname{\mathsf{P}}_{\theta}\subset\operatorname{\mathsf{G}}$ that normalizes $\mathfrak{u}_{\theta}$, and the $\theta$-flag manifold is the compact manifold $\operatorname{\mathcal{F}}_{\theta}:=\operatorname{\mathsf{G}}/\operatorname{\mathsf{P}}_{\theta}$. Note that the left action of $\operatorname{\mathsf{G}}$ on $\operatorname{\mathcal{F}}_{\theta}$ endows $\operatorname{\mathcal{F}}_{\theta}$ with the structure of a homogeneous $\operatorname{\mathsf{G}}$-space.
The (closed) positive Weyl chamber is
$\mathfrak{a}_{\operatorname{\mathsf{G}}}^{+}:=\{X\in\mathfrak{a}:\alpha(X)\geq 0\text{ for all }\alpha\in\Delta_{\operatorname{\mathsf{G}}}\}.$
The $\mathsf{KAK}$-decomposition theorem states that every $g\in\operatorname{\mathsf{G}}$ can be written as $g=k_{1}e^{\kappa(g)}k_{2}$ for unique $\kappa(g)\in\mathfrak{a}_{\operatorname{\mathsf{G}}}^{+}$ and some $k_{1}$ and $k_{2}$ in $\mathsf{K}$, which is the maximal compact subgroup of $\operatorname{\mathsf{G}}$ whose Lie algebra is $\mathfrak{k}$. This defines for us the Cartan projection
$\kappa:\operatorname{\mathsf{G}}\to\mathfrak{a}_{\operatorname{\mathsf{G}}}^{+}.$
In the case when $\operatorname{\mathsf{G}}=\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, we may choose $\mathfrak{k}$ to be $\mathfrak{su}(d,\operatorname{\mathbb{K}})$, which is the set of traceless, skew-Hermitian, $d\times d$, $\operatorname{\mathbb{K}}$-valued matrices, and choose $\mathfrak{a}_{\operatorname{\mathsf{G}}}$ and $\mathfrak{a}_{\operatorname{\mathsf{G}}}^{+}$ to be $\mathfrak{a}$ and $\mathfrak{a}^{+}$ as defined in Section 2.1. One can then verify that the restricted simple roots defined here agree with the ones defined in Section 2.1, and that $\Delta_{\operatorname{\mathsf{G}}}=\{\alpha_{k}:k\in\Delta\}$. Also, the Cartan projection defined here agrees with the one defined in Section 2.1.
Let $\theta\subset\Delta_{\operatorname{\mathsf{G}}}$ be a non-empty, symmetric subset. A discrete subgroup $\Gamma\subset\operatorname{\mathsf{G}}$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov if it is finitely generated, and there exists $a,C>0$ such that
$\alpha_{k}(\kappa(\gamma))\geq a|\gamma|-C$
for all $\gamma\in\Gamma$, and $k\in\theta$, where $|\gamma|$ is the word length $\gamma$ with respect to some fixed finite generating set for $\Gamma$. By Theorem 2.2, in the case when $\operatorname{\mathsf{G}}=\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, this agrees with the notion of $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroups defined in Section 2.3. In general, every $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup $\Gamma\subset\operatorname{\mathsf{G}}$ admits a unique non-empty, closed, $\Gamma$-invariant, subset $\Lambda_{\theta}(\Gamma)\subset\operatorname{\mathcal{F}}_{\theta}$ whose points are pairwise transverse, and on which the $\Gamma$-action is strongly dynamics preserving, i.e. if $\{\gamma_{n}\}$ is an infinite sequence in $\Gamma$, then up to taking a subsequence, there exist $a,b\in\Lambda_{\theta}(\Gamma)$ such that the sequence $\{\gamma_{n}\}$, when restricted to the set of flags in $\operatorname{\mathcal{F}}_{\theta}$ that are transverse to $b$, converges uniformly on compact sets to the constant map whose image is $a$. We refer to $\Lambda_{\theta}(\Gamma)$ as the $\theta$-limit set of $\Gamma$.
Results in [20] imply the following.
@@ 4.1. Anosov groups in general semisimple Lie groups
Suppose that $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a $\mathsf{P}_{k}$-Anosov subgroup that is strongly irreducible, let $\mathsf{G}$ be the real Zariski closure of $\Gamma$ in $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, and let $\operatorname{\mathsf{G}}^{0}$ denote the identity component of $\operatorname{\mathsf{G}}$. Then:
  1. (1)
    $\operatorname{\mathsf{G}}^{0}$ has trivial center and is semisimple.
  2. (2)
    There exists a non-empty $\theta\subset\Delta_{\operatorname{\mathsf{G}}}$ such that $\Gamma$ is $\mathsf{P}_{\theta}$-Anosov, and for which there is a $\mathsf{G}$-equivariant smooth embedding $\iota:\mathsf{G}/\mathsf{P}_{\theta}\hookrightarrow\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$ such that the $\theta$-limit set $\Lambda_{\theta}(\Gamma)\subset\mathsf{G}/\mathsf{P}_{\theta}$ of $\Gamma$ in satisfies
    $\iota(\Lambda_{\theta}(\Gamma))=\Lambda_{k}(\Gamma).$
Proof.
Let $\Gamma^{\prime}$ and $\operatorname{\mathsf{G}}^{\prime}$ respectively denote the preimages of $\Gamma$ and $\operatorname{\mathsf{G}}$ under the quotient map $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})\to\mathsf{PSL}(d,\operatorname{\mathbb{K}})$. Observe that $\operatorname{\mathsf{G}}^{\prime}$ is the real Zariski closure of $\Gamma^{\prime}$.
First, we prove (1). Since $\Gamma$ is strongly irreducible, it follows that the identity component $(\operatorname{\mathsf{G}}^{\prime})^{0}$ acts irreducibly on $\operatorname{\mathbb{K}}^{d}$, and hence is reductive. We may then apply Schur’s lemma to deduce that the center of $(\operatorname{\mathsf{G}}^{\prime})^{0}$ acts on $\operatorname{\mathbb{K}}^{d}$ by scaling, and hence lies in the center of $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})$. In particular, $(\operatorname{\mathsf{G}}^{\prime})^{0}$ is semisimple. It now follows that $\operatorname{\mathsf{G}}^{0}=\pi((\operatorname{\mathsf{G}}^{\prime})^{0})$ is semisimple and has trivial center.
To prove (2), notice that it suffices to prove the statement for $\operatorname{\mathsf{G}}^{\prime}$ and $\Gamma^{\prime}$ in place of $\operatorname{\mathsf{G}}$ and $\Gamma$. Let $\eta=\eta_{k}:\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})\to\operatorname{\mathsf{SL}}(\wedge^{k}\operatorname{\mathbb{K}}^{d})$ be the representation induced by the natural linear action of $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})$ on $\wedge^{k}\operatorname{\mathbb{K}}^{d}$, and let $\chi:\mathfrak{a}_{\operatorname{\mathsf{G}}^{\prime}}\to\operatorname{\mathbb{R}}$ be the highest weight of $\eta|_{\operatorname{\mathsf{G}}^{\prime}}$. We may write
$\chi=\sum_{\alpha\in\Delta_{\operatorname{\mathsf{G}}^{\prime}}}c_{\alpha}\omega_{\alpha},$
where $\omega_{\alpha}$ is the restricted simple weight corresponding to $\alpha$ and each $c_{\alpha}$ is a non-negative integer. Let $\theta:=\{\alpha\in\Delta:c_{\alpha}>0\}$. Since $\Gamma$ is $\operatorname{\mathsf{P}}_{k}$-Anosov, $\eta(\Gamma)$ is $\operatorname{\mathsf{P}}_{1}$-Anosov (see Item (VII) in Appendix 2.6), and so contains a proximal element. In the language of Guichard–Gueritaud–Kassel–Wienhard (see [20, Lemma 3.2]), we have just observed that $\eta|_{\operatorname{\mathsf{G}}^{\prime}}$ is $\theta$-proximal, so we may apply [20, Propositions 3.3 and 3.5] to deduce (2). ∎
Given Lemma 4.3 part (1), to prove Theorem 4.1, it now suffices to show that the real Zariski closure of $\operatorname{\mathsf{G}}$ in $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is simple. The proof of this is divided into two steps which are written as Lemmas 4.5 and 4.9 below. The former shows that the real Zariski closure of $\Gamma$ has no compact factors, while the latter shows that it has at most one simple factor.
4.2. Ruling out compact simple factors
We prove some general results which show that if $\Gamma$ satisfies the hypothesis Theorem 4.1, then the connected component of the identity of the real Zariski closure of $\Gamma$ has no compact simple factors.
We start by recording a standard fact about irreducible representations of products of semisimple Lie groups.
@@ 4.2. Ruling out compact simple factors
Suppose $\mathsf{H}_{1},\mathsf{H}_{2}\subset\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})$ are commuting connected reductive Lie groups and $\mathsf{H}:=\mathsf{H}_{1}\mathsf{H}_{2}$ acts irreducibly on $\operatorname{\mathbb{K}}^{d}$. Then there exist irreducible representations $\rho_{j}:\mathsf{H}_{j}\rightarrow\operatorname{\mathsf{SL}}(d_{j},\operatorname{\mathbb{K}})$ and a linear identification
@@
Since $\mathsf{H}_{1}$ is reductive, we may decompose its action on $\operatorname{\mathbb{K}}^{d}$ into its irreducible factors, i.e. there exist $m_{1},\dots,m_{k},n_{1},\dots,n_{k}\in\operatorname{\mathbb{N}}$ and irreducible representations $\tau_{j}:\mathsf{H}_{1}\rightarrow\operatorname{\mathsf{SL}}(m_{j},\operatorname{\mathbb{K}})$ so that we can identify
@@
and $\mathsf{H}_{1}$ acts by $(\tau_{1}^{n_{1}},\dots,\tau_{k}^{n_{k}})$. Since $\mathsf{H}_{1}$ and $\mathsf{H}_{2}$ commute, $\mathsf{H}_{2}$ preserves each $(\operatorname{\mathbb{K}}^{m_{j}})^{n_{j}}\subset\operatorname{\mathbb{K}}^{d}$. Then, since $\mathsf{H}$ acts irreducibly, we must have $k=1$.

### v2

@@ 4. Rigidity
$\delta^{\alpha_{k}}(\Gamma)=2,$
In the case when the Zariski closure of $\Gamma$ is simple, we originally had a proof of Theorem 4.1 that was more dynamical in flavor (a sketch of this argument is provided at the end of the section). We thank Andres Sambarino for suggesting that we use Chow’s theorem [13] instead.
@@
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex subgroup and $\partial\Gamma$ is a two-sphere, then $\Lambda_{k-1,k,k+1}(\Gamma)$ is a complex submanifold of $\operatorname{\mathcal{F}}_{k-1,k,k+1}(\operatorname{\mathbb{C}}^{d})$ (where we use the convention $\operatorname{\mathcal{F}}_{0,1,2}(\operatorname{\mathbb{C}}^{d})=\operatorname{\mathcal{F}}_{1,2}(\operatorname{\mathbb{C}}^{d})$ and $\operatorname{\mathcal{F}}_{d-2,d-1,d}(\operatorname{\mathbb{C}}^{d})=\operatorname{\mathcal{F}}_{d-2,d-1}(\operatorname{\mathbb{C}}^{d})$).
Assuming Lemma 4.2, we will now prove Theorem 4.1. Recall that by Schur’s lemma, any irreducible Lie subgroup of $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{C}})$ is semisimple, and its center lies in the center of $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{C}})$. Therefore, any irreducible Lie subgroup of $\mathsf{PSL}(d,\operatorname{\mathbb{C}})$ is semisimple with trivial center.
We use the following standard fact about irreducible representations of products of semisimple Lie groups.
@@
Suppose $\mathsf{H}_{1},\mathsf{H}_{2}\subset\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})$ are commuting connected semisimple Lie groups and $\mathsf{H}:=\mathsf{H}_{1}\mathsf{H}_{2}$ acts irreducibly on $\operatorname{\mathbb{K}}^{d}$. Then there exist irreducible representations $\rho_{j}:\mathsf{H}_{j}\rightarrow\operatorname{\mathsf{SL}}(d_{j},\operatorname{\mathbb{K}})$ and a linear identification
@@
Since $\mathsf{H}_{1}$ is semisimple, we may decompose its action on $\operatorname{\mathbb{K}}^{d}$ into its irreducible factors, i.e. there exist $m_{1},\dots,m_{k},n_{1},\dots,n_{k}\in\operatorname{\mathbb{N}}$ and irreducible representations $\tau_{j}:\mathsf{H}_{1}\rightarrow\operatorname{\mathsf{SL}}(m_{j},\operatorname{\mathbb{K}})$ so that we can identify
@@
and $\mathsf{H}_{1}$ acts by $(\tau_{1}^{n_{1}},\dots,\tau_{k}^{n_{k}})$. (See, for example, Serre [30, Chap. VI.3].) Since $\mathsf{H}_{1}$ and $\mathsf{H}_{2}$ commute, $\mathsf{H}_{2}$ preserves each $(\operatorname{\mathbb{K}}^{m_{j}})^{n_{j}}\subset\operatorname{\mathbb{K}}^{d}$. Then, since $\mathsf{H}$ acts irreducibly, we must have $k=1$.